# Neuron Empirical Gradient: Discovering and Quantifying Neurons’ Global Linear Controllability

### v2

@@ 4.3 Knowledge Attribution Evaluation
Finally, we evaluate NeurGrad’s ability to locate important neurons. Specifically, for 1000 prompts, we identify the top-$K$ neurons ($K=1,4,16$) with the highest attribution scores using CG, IG, LPI, and NeurGrad. We enhance neuron activations by shifting positive neurons from 0.1 to 1 and negative neurons from -0.1 to -1, both in increments of 0.1. Figure 3 reports the output shifts of target tokens on Llama2-7b, with different activation shifts and top-k neurons selected from different attribution methods. Figure 3 demonstrates that NeurGrad outperforms baselines in attribution accuracy. This superiority stems from NeurGrad’s precise measurement of NEG and its inclusion of negative neurons in intervention, unlike IG and LPI, which only consider positive neurons despite their equal ratio to negative neurons (Table 2). See details in § C.
@@ 5 Understanding Neurons’ Controllability
This section deepens our understanding of neuron controllability: the ability to precisely adjust PLM output probabilities by modifying neuron activations, using NEGs estimated by NeurGrad.
BERTbase/large/wwm Llama2-7/13/70B
Pos. ratio .5019/.5008/.4996 .4604/.4664/.4484
Neg. ratio .4981/.4992/.5004 .4592/.4660/.4480
Table 2: The pos/neg neuron ratios over 1000 prompts.
@@ 5.1 How Are NEGs Distributed?
Do neurons have polarity preference? Table 2 reports the ratios of positive/negative neurons across 1000 prompts in the six PLMs, showing that the numbers of positive/negative neurons are nearly equivalent. It indicates that PLMs show no preference for their neurons’ polarities, suggesting that enhancing or suppressing neurons should be guided by their gradient polarity rather than merely increasing or decreasing their activations Dai et al. (2022). See detailed distribution analysis in § B.2,B.3.
Does only a few neurons exhibit strong gradients? Figure 4 shows the cumulative distribution of NEG for all neurons in PLMs. Rising curves are steady and do not converge until all neurons are present, indicating that most neurons can affect the model’s output probabilities.
@@ 5.2 Does Linearity Hold for Multi-neuron?
We conduct multi-neuron intervention experiments to investigate: can output shifts be predicted when modifying multiple neuron activations? We randomly sample $N$ neurons and enhance them by shifting their activations according to their polarities measured by NeurGrad, in which positive and negative activation shifts are applied to positive/negative neurons. We experiment on BERTbase and Llama2-7B using neuron sizes of $2^{N}$ ($0\leq N\leq 12$) across 1000 prompts.
Figure 5 shows the average Corr across all prompt-neuron pairs with an enhancement range of [0, 0.5] and a step size of 0.01. We observe that even Corr decreases as more neurons are involved due to neuron interactions, a strong correlation ($\geq$0.7) remains even with 4096 neurons in both PLMs. Additionally, Figure 5 (right y-axis) shows that involving more neurons can cause larger output shifts, suggesting that the output shifts caused by individual neurons can be accumulated. Note that despite the small average output shifts in Llama2 due to the small NEG magnitudes per neuron (see § B.2 for details), significant probability changes can still be observed when more than 1024 neurons are involved in Llama2-7B models with specific neuron combinations. Moreover, our analysis with larger enhancement ranges reveals consistent trends: increasing the shift range reduces Corr, making output shifts less predictable. See § D for experiments with different ranges.
The analysis above demonstrates that modifying neuron activations, as guided by NeurGrad, enables partial prediction of output shifts. However, the number of modified neurons and the range of modifications require careful examination.
5.3 Local Linear Approximation Hypothesis
We present the Local Linearity Approximation Hypothesis to explain neuron linearity, based on three key observations: (1) expanding the shift range reduces Corr (Figure 2); (2) PLMs with more parameters diminish the importance of individual neurons, leading to higher Corr (Figure 2); (3) involving more neurons weakens linearity (Figure 5).
Let $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a PLM, where $x\in\mathbb{R}^{n}$ represents neuron activations and $f(x)$ represents the output token probabilities. Within a constrained region, the influence of a single neuron $x_{i}$ can be locally approximated as: $f(x_{i}+\delta e)\approx f(x_{i})+\frac{\partial f}{\partial x_{i}}\delta,$ for small $\delta$, where $e_{i}$ is the unit basis vector. This follows from the first-order Taylor expansion, reflecting the model’s local differentiability with respect to neuron activations.
@@ 6 Skill Neuron Probing using NeurGard
While NEG and NeurGrad enable neuron-level modifications to model outputs, the variability of neurons’ NEG values across different prompts limits their ability to support modifications for specific types of knowledge, such as language skills. These skills often involve handling a range of prompts that require linguistic diversity. This section explores whether NEG can effectively capture general language skills linked to diverse prompts through skill neuron probing. Skill neuron probing seeks to identify neurons that encode the ability to solve language tasks. Note that prior work (Wang et al., 2022; Song et al., 2024) explored the effectiveness of using neuron activations for this purpose, yet we focus on neurons’ NEG.
@@ 6.1 Task Definition
Following Wang et al. (2022), we formulate the skill neuron probing task as follows. A dataset conveying specific language skills $\mathcal{D}$ consists of language sequence pairs, including knowledge inquiries $\mathcal{Q}=\{q_{1},...,q_{|\mathcal{T}|}\}$ and answer sequences $\mathcal{A}=\{a_{1},...,a_{|\mathcal{T}|}\}$, where arbitrary $a_{i}$ belongs to the answer candidate set $\mathcal{\hat{A}}_{\text{cands}}$. For example, in the sentiment classification task, $Q$ is the documents set, and $A$ is the ground-truth sentiment labels. We then build classifiers that take behaviors of arbitrary neuron subset $\mathcal{N}_{s}\subseteq\mathcal{N}$ as features to indicate the correct answer sequences $a_{i}$ for the knowledge inquiry $q_{i}$. $\mathcal{N}$ refers to all the neurons. [^9] We focus on intermediate outputs (neurons) of FF layers.
Our skill neuron prober aims to find $\mathcal{N}_{s}^{*}$ that can achieve optimal accuracy over the target dataset $\mathcal{D}$.
@@
Here, $f(\mathcal{N}_{s},q_{i})$ is the output of the classifier $F$ using the neuron subset $\mathcal{N}_{s}$ for the prompt $q_{i}$. $\mathbbm{1}[X=Y]$ is an indicator function that equals $1$ if $X$ matches $Y$, and 00 otherwise.
@@ 6.2 Evaluation Benchmark: MCEval8K
This section, we create a multi-choice knowledge evaluation benchmark-MCEval8K for skill neuron probing. As skill neuron probing requires a fixed target token, previous studies (Wang et al., 2022; Song et al., 2024) relied on multi-choice datasets that forces PLMs to generate a single-token option label (A, B, etc.) named for category labels (A: positive, B: negative, etc.). While we adopt a similar setup, earlier work focused on small PLMs with datasets limited to basic language understanding tasks, which are insufficient for evaluating LLMs.
@@
To address this, we introduce MCEval8K, a diverse multi-choice benchmark covering 22 tasks across six genres, designed to assess a wide range of language skills, incorporating most datasets from previous studies. Since tasks vary in different sizes, with some, such as cLang-8 (Rothe et al., 2021; Mizumoto et al., 2011), containing millions of data points, we standardize the evaluation by limiting each task to 8K queries. [^10] Only the Stereoset task has fewer than 8K queries due to the limited size of the original dataset. It minimizes unnecessary computational costs while ensuring consistency across tasks. We also ensure the number of ground-truth options per task is balanced to eliminate bias introduced by imbalanced classification. The skill genres, tasks, and datasets information are shown below (detailed in § F).
Linguistic: Part-of-speech tagging (POS), phrase-chunking (CHUNK), named entity recognition (NER), and grammatical error detection (GED). Content classification: Sentiment (IMDB), topic classification (Agnews), and Amazon reviews with numerical labels (Amazon). Natural language inference (NLI): textual entailment (MNLI), paraphrase identification (PAWS), and grounded commonsense inference (SWAG). Factuality: Fact-checking (FEVER), factual knowledge probing (MyriadLAMA), commonsense knowledge (CSQA), and temporary facts probing (TempLAMA). Self-reflection: Examine PLMs’ internal status, including hallucination (HaluEval), toxicity (Toxic), and stereotype (Stereoset) detections. Multilinguality: We select tasks with multilingual queries, including language identification (LTI), multilingual POS-tagging on Universal Dependencies (M-POS), Amazon review classification (M-Amazon), factual knowledge probing (mLAMA) and textual entailment (XNLI).
@@
Figure 6: MCEval8K accuracies on Llama2-7B across tasks in zero-shot and few-shot settings, reported for Rand (random guess), TProb (token probability), and two proposed probers. Legends show average accuracies.
7 NEG as Knowledge Feature
We train skill neuron probers based on NeurGrad’s estimated gradients to investigate whether and how NEG encodes language knowledge.
7.1 Gradient-based Skill Neuron Prober
For each task dataset $\mathcal{D}$, we split it into: training set $\mathcal{D}_{\text{train}}$ to train the classifiers, validation set $\mathcal{D}_{\text{valid}}$ to decide hyperparameters, and test set $\mathcal{D}_{\text{test}}$ for evaluation, with the ratio of 6:1:1. We train three probers with different designs for comparison.
@@ 7.1 Gradient-based Skill Neuron Prober
Polarity-based majority vote (Polar-prober) adopts a simple majority-vote classifier, taking each neuron in $\mathcal{N}_{s}$ as one voter. A polarity-based classifier leverages the polarity of neurons (positive or negative) as features for classification. Given $\mathcal{D}_{\text{train}}=\{(q_{i},a_{i})\}$ and any neuron $n_{k}\in\mathcal{N}$, we identify the polarity as feature $\mathbf{x}_{q_{i},a_{i}}^{n_{k}}$ for each $(q_{i},a_{i})$ pair. For each $n_{k}$, we calculate the ratio of being positive and negative across all $|\mathcal{D}_{\text{train}}|$ examples and the dominant polarity is identified as their global polarity $\bar{\mathbf{x}}^{n_{k}}$. Neurons with more consistent polarity are ranked higher.
To make prediction of $q_{i}$, we measure all polarities of $\mathbf{x}_{q_{i},a_{j}}^{n_{k}}$, where $a_{j}\in\hat{\mathcal{A}}_{\text{cands}},n_{j}\in\mathcal{N}_{s}^{*}$. The prediction of each $q_{i}$ is made as follows:
$f(\mathcal{N}_{s}^{*},q_{i})$ $=\underset{a_{j}\in\hat{\mathcal{A}}_{\text{cands}}}{\arg\max}\sum_{n_{k}\in\mathcal{N}_{s}^{*}}\mathbbm{1}[\mathbf{x}_{q_{i},a_{j}}^{n_{k}}=\bar{\mathbf{x}}^{n_{k}}]$ (4)
We identify the optimal size of $\mathcal{N}_{s}^{*}$ with $\mathcal{D}_{\text{valid}}$.
Magnitude-based majority vote (Magn-prober) utilizes gradient magnitudes as features for a majority-vote classifier. During training, for a specific $q_{i}$ and $n_{k}$, we compare the gradients between $a\in\hat{\mathcal{A}}_{\text{cands}}$. Neurons that consistently exhibit the largest or smallest gradients for the ground truth $a_{i}$ compared to other candidates are used as skill indicators. We record each neuron’s preference for being either the largest or smallest. Neurons exhibiting more consistent behavior are assigned higher importance and identified as skill neurons. During inference, similar to Eq. 4, the prediction is made by selecting $a_{j}$ that satisfies the majority of $n_{k}\in\mathcal{N}_{s}^{*}$. This prober is designed to compare against the polarity-based prober, aiming to examine whether NEGs’ magnitude can bring additional skill information compared to polarity.
@@ 7.2 Experimental Setup
Dataset & Prompt Since our probing method restricts the output sequence length to one, we carefully craft instructions and options for all datasets in MCEval8K through human effort. We evaluate both zero-shot and few-shot settings, ensuring in few-shot experiments that all candidate tokens appear once in the demonstrations to prevent majority label bias Zhao et al. (2021). See § J for the designed instructions for all tasks.
Prober During validation, we select the optimal neuron size for majority-vote probers from $2^{n}(0\leq n\leq 13)$. See § G for detailed prober settings.
Model We perform skill neuron probing on Llama2-7B using three probers, all datasets in MCEval8K, and the full training set (6000) per task. For Llama2-70B, due to high cost, we probe one dataset per genre—NER, Agnews, PAWS, CSQA, HaluEval, and mLAMA—using 1024 training examples and only train major-vote probers.
@@ 7.3 Result and Analysis
Skill neuron-based classifier accuracy is compared to two baselines: random guessing (Rand), and answer token probability-based classification (LM-Prob) which selects the candidate token with the highest probability as the prediction, serving as a benchmark for the LLMs’ prompting performance.
Empirical gradients encode language skills. Figure 6 shows accuracies for all tasks in MCEval8K using the Llama2-7B, with both zero- and few-shot settings. The results demonstrate that LM-Prob outperforms Rand, indicating that Llama2-7B is capable of understanding instructions and recalling skills from its parameters. We also confirm the effectiveness of our skill neuron probers in addressing language tasks. The two simple major-vote classifiers outperform LM-Prob in both settings. The per-task classification accuracies in Figure 6 show that skill neurons effectively represent diverse language skills, achieving consistently high results across tasks. See Table 7,8 for accuracy values.
Tasks Llama2-7B Llama2-70B
LM-Prob Magn-Prober LM-Prob Magn-Prober
NER .3610 .4980 .7900 .8170
Agnews .5880 .7020 .7630 .8240
PAWS .5240 .8150 .7790 .8460
CSQA .6100 .6390 .7540 .7630
HaluEval .5200 .7830 .7530 .8250
mLAMA .6080 .6370 .7430 .7600
Table 3: Accuracies of 6 tasks on Llama2-7B and -70B.
Larger PLMs excel in skill recall. Table 3 compares LM-Prob and Magn-prober across six tasks in the few-shot setting between Llama-7B and -70B. Llama-70B outperforms Llama-7B in both LM-Prob and skill neuron probing. However, the difference between LM-Prob and Magn-prober is smaller in Llama2-70B than in Llama2-7B, indicating the large model’s strong ability to recall knowledge from its parameters. See § H for further analysis on the properties of skill neurons: efficiency, generality, and inclusivity.

### v3

@@ 3.3 Knowledge Attribution Evaluation
We further evaluate NeurGrad’s ability to identify important neurons. For 1000 prompts, we select the top-$K$ neurons ($K=1,2^{2},2^{4}$) using CG, IG, LPI, and NeurGrad values, and then enhance their activations by increasing activations of positive neurons and decreasing those of negative neurons. The activation shift is conducted within the range of [0.1, 1] with a step size of 0.1. Figure 3 shows output shifts on Llama2-7B under these interventions. NeurGrad consistently outperforms baselines, due to its accurate NEG estimation and inclusion of both positive and negative neurons, unlike IG and LPI, which only consider positive ones, despite their equal distribution (Table 1). See § C for details.
@@ 4 Understanding Neurons’ Controllability
This section explores neuron controllability: the ability to precisely adjust PLM output probabilities by modifying neuron activations. We use NEGs estimated by NeurGrad to achieve this.
@@ 4.1 How Are NEGs Distributed?
Do only a few neurons exhibit strong gradients?
Figure 4 presents the cumulative NEG distribution for all neurons, showing steady growth that only converges when all neurons are included. This indicates that most neurons influence the model’s output probabilities.
Do neurons have polarity preference?
Table 1 (§ 2) shows the ratios of positive and negative neurons [^9] We used ground-truth NEGs here to avoid flipping the polarity of neurons with NEGs near zero. across 1000 prompts, each with 100 randomly sampled neurons in seven PLMs, revealing nearly equal numbers. This suggests that PLMs have no polarity preference; interventions should consider gradient polarity rather than simply increasing or decreasing their activations Dai et al. (2022). See more detailed analysis on NEGs’ distribution in § B.2 and § B.3.
@@ 4.2 Does Linearity Hold for Multi-neuron?
We explore whether output shifts can be predicted when intervening multiple neurons. For each of 1000 prompts, we randomly sample $N$ neurons and shift their activations on the basis of polarity measured by NeurGrad, applying positive/negative shifts to positive/negative neurons. We experiment on BERTbase and Llama2-7B using neuron sizes of $2^{N}$ ($0\leq N\leq 12$).
Figure 5 shows the average correlation ($r$) between predicted and actual output shifts across all prompt-neuron pairs, using an enhancement range of [0, 0.5] and a 0.01 step size. While $r$ decreases as more neurons are involved due to neuron interactions, it remains strong ($\geq$0.7) even with $2^{12}$ neurons in both PLMs. The figure also shows that larger neuron sets cause greater output shifts, suggesting an additive effect. Although Llama2 shows small average output shifts (due to the lower per-neuron NEG; see § B.2), significant changes still occur when over $2^{10}$ neurons are intervened. Experiments with larger enhancement ranges show similar trends: greater shifts load to lower $r$, reducing predictability (see § D for experiments with different ranges).
The analysis above shows that NeurGrad enables partial prediction of output shifts. However, both the number of modified neurons and the modification range require careful consideration.
4.3 Why Do Neurons Exhibit Linearity?
We propose the local linearity approximation hypothesis to explain neuron linearity, based on three observations: i) larger shift ranges reduce $r$, ii) larger PLMs show higher $r$ due to the reduced influence of individual neurons (Figure 2), and iii) involving more neurons weakens linearity (Figure 5). This aligns with the first-order Taylor expansion under local differentiability.
Local Linear Approximation Hypothesis Let $\mathbf{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ denote the mapping from a model’s FF activation layer $\mathbf{x}\in\mathbb{R}^{n}$ to its output probability vector $\mathbf{f}(\mathbf{x})$. We have: $f_{j}(\mathbf{x}^{\prime})$ $=f_{j}(\mathbf{x})+\frac{\partial f_{j}}{\partial x_{i}}(\mathbf{x})\,\Delta x_{i}+o(\Delta x_{i})$ if we consider a perturbation $\Delta x_{i}$ on $i$-th neuron, with perturbed vector $\mathbf{x}^{\prime}=\mathbf{x}+\Delta x_{i}\mathbf{e}_{i}$. Here $(\mathbf{e}_{i})_{j}=\delta_{ij}$ is a unit vector. Then, we define $\frac{\partial f_{j}}{\partial x_{i}}(\mathbf{x})$ as NEG.
@@ 5 Skill Neuron Probing using NeurGard
While NEG and NeurGrad allow neuron-level output control, variation in NEG across prompts limits their use for modifying specific language skills, which require handling diverse inputs. This section investigates whether NEG can capture general language skills through skill neuron probing, which identifies neurons linked to task-solving capabilities. Unlike prior work (Wang et al., 2022; Song et al., 2024) focusing on neuron activations, we utilize NEG for finding skill neurons.
@@ 5.1 Task Definition
Following Wang et al. (2022), we define skill neuron probing as follows. A skill dataset $\mathcal{D}$ consists of language sequence pairs: knowledge inquiries $\mathcal{Q}=\{q_{1},...,q_{|\mathcal{T}|}\}$ and corresponding answers $\mathcal{A}=\{a_{1},...,a_{|\mathcal{T}|}\}$, where each $a_{i}$ is from a candidate set $\mathcal{\hat{A}}_{\text{cands}}$. For example, in sentiment classification, $Q$ is the set of documents and $A$ the sentiment labels. We train classifiers using the behavior of a neuron subset $\mathcal{N}_{s}\subseteq\mathcal{N}$ as features to predict the correct answer $a_{i}$ for each $q_{i}$, where $\mathcal{N}$ is the full neuron set.  [^10] We focus on intermediate outputs (neurons) of FF layers.
Our skill neuron probe finds the optimal neuron subset $\mathcal{N}_{s}^{*}$ that maximizes accuracy on the dataset $\mathcal{D}$.
@@
Here, $f(\mathcal{N}_{s},q_{i})$ is the classifer’s prediction using neuron subset $\mathcal{N}_{s}$ for input $q_{i}$ and $\mathbbm{1}[X=Y]$ is $1$ if $X=Y$, otherise 00.
@@ 5.2 Evaluation Benchmark: MCEval8K
In this section, we introduce a benchmark dataset for skill neuron probing. Since probing requires a fixed target token, previous work (Wang et al., 2022; Song et al., 2024) used on multiple-choice datasets with single-token labels (e.g., A: positive, B: negative). While we follow a similar setup, earlier studies focused on small PLMs and probed them on basic tasks, which are insufficient for evaluating LLMs.
@@
To address this limitation, we create MCEval8K, a diverse multiple-choice benchmark covering 22 language understanding tasks across six skill genres, incorporating most datasets from previous studies. To ensure consistency and reduce computational cost, we cap each task at 8K queries, [^11] Only the Stereoset task has fewer than 8K queries due to the limited size of the original dataset. even for large datasets like cLang-8 (Mizumoto et al., 2011; Rothe et al., 2021). We also balance the number of correct options per task to avoid classification bias. Task and dataset details are provided below and in § F.
Linguistic: Part-of-speech tagging (POS), text chunking (CHUNK), named entity recognition (NER), and grammatical error detection (GED).
Content classification: Sentiment (IMDB), topic (Agnews), and Amazon reviews with numerical labels (Amazon).
Natural language inference (NLI): Textual entailment (MNLI), paraphrase identification (PAWS), and grounded commonsense inference (SWAG).
Factuality: Fact-checking (FEVER), knowledge probing (MyriadLAMA), commonsense QA (CSQA), and temporal fact probing (TempLAMA).
Self-reflection: Hallucination (HaluEval), toxicity (Toxic), and stereotype detections (Stereoset).
Multilinguality: Language identification (LTI), multilingual POS tagging (M-POS), sentiment classification (M-Amazon), knowledge probing (mLAMA), and textual entailment (XNLI).
@@
Figure 6: MCEval8K accuracies on Llama2-7B across tasks in zero-shot and few-shot settings, reported for Rand (random guess), LM-Prob (token probability), and three proposed probes. Legends show average accuracies.
Model NER Agnews PAWS CSQA HaluEval mLAMA
LM Act Mag LM Act Mag LM Act Mag LM Act Mag LM Act Mag LM Act Mag
Llama3.2-3B .598 .633 .618 .757 .857 .776 .500 .843 .831 .551 .681 .691 .650 .822 .780 .473 .563 .570
Llama2-7B .361 .453 .498 .588 .849 .702 .524 .825 .815 .610 .613 .639 .520 .788 .783 .608 .622 .637
Qwen2.5-7B .871 .877 .877 .751 .858 .755 .574 .889 .873 .768 .816 .814 .637 .821 .773 .784 .795 .785
Llama3.1-8B .815 .826 .833 .770 .883 .812 .500 .860 .854 .674 .733 .740 .674 .807 .798 .762 .781 .785
Llama2-70B .790 - .817 .763 - .824 .779 - .846 .754 - .763 .753 - .825 .743 - .760
Table 3: Accuracy of six tasks across five LLMs: LM-Prob, Activation-based, and Magn-Probe.
5.3 NEG-based Skill Neuron Probe
We train skill neuron probes using NeurGrad’s estimated NEG to examine how NEG encodes general language skills. Each skill dataset $\mathcal{D}$ is split into: training, validation, and test sets ($\mathcal{D}_{\mathrm{train}}$, $\mathcal{D}_{\mathrm{valid}}$, and $\mathcal{D}_{\mathrm{test}}$), with a ratio of 6:1:1. We compare the following three probes with different designs.
@@ 5.3 NEG-based Skill Neuron Probe
Polar-Probe is a majority-vote classifier, where each neuron votes based on its polarity (positive or negative). For each training pair $(q_{i},a_{i})$ in $\mathcal{D}_{\mathrm{train}}$ and neuron $n_{k}$, we record its polarity as feature $\bm{x}_{q_{i},a_{i}}^{n_{k}}$. The dominant polarity across $\mathcal{D}_{\mathrm{train}}$ defines the neuron’s global polarity $\bar{\bm{x}}^{n_{k}}$, and neurons with more consistent polarity are ranked higher.
To predict $q_{i}$, we compare polarities $\bm{x}_{q_{i},a_{j}}^{n_{k}}$ for each candidate $a_{j}\in\hat{\mathcal{A}}_{\text{cands}}$ and $n_{j}\in\mathcal{N}_{s}^{*}$:
$f(\mathcal{N}_{s}^{*},q_{i})$ $=\underset{a_{j}\in\hat{\mathcal{A}}_{\text{cands}}}{\arg\max}\sum_{n_{k}\in\mathcal{N}_{s}^{*}}\mathbbm{1}[\bm{x}_{q_{i},a_{j}}^{n_{k}}=\bar{\bm{x}}^{n_{k}}]$ (4)
The optimal size of $\mathcal{N}_{s}^{*}$ is selected using $\mathcal{D}_{\text{valid}}$.
Magn-Probe uses NEG magnitudes as features for a majority-vote classifier. For each training pair $(q_{i},a_{i})$ in $\mathcal{D}_{\mathrm{train}}$ and neuron $n_{k}$, we compare NEGs across $a\in\hat{\mathcal{A}}_{\text{cands}}$. Neurons that consistently show the highest or lowest NEGs for the correct answer $a_{i}$ are marked as skill indicators, along with their preference for being highest or lowest. More consistent neurons are ranked higher. At inference, predictions follow the same voting rule as in Eq. 4. This probe evaluates whether NEG magnitudes can encode skill information.
Tree-Probe is designed to analyze the impact of accounting for interdependencies among skill neurons. We use the index (non-negative integers) of $a\in\hat{\mathcal{A}}_{\text{cands}}$ with the most significant NEGs as features for training a random forest classifier. The hyperparameters include the number of trees (#n_trees) and layers (#n_layers) used in each tree. See more details in § G.2.
@@ 5.4 Experimental Setup
Dataset & Prompt. We split each MCEval8K dataset into train, validation, and test by the ratio of 6:1:1, ensuring balanced correct-answer tokens across subsets to avoid majority label bias Zhao et al. (2021). We manually craft instructions and options for all MCEval8K tasks to ensure single-token outputs, considering both zero-shot and few-shot settings. See § I for all task instructions.
Probe.
For majority-based probes, we select the optimal neuron size from $2^{n}(0\leq n\leq 13)$ using $\mathcal{D}_{\mathrm{valid}}$. For the Tree-Probe (random forest), we report accuracy using scikit-learn’s default settings, where the optimal subset of features is selected automatically: 100 trees with no depth limitation. See § G.1G.2 for details.
Models.
We probe Llama2-7B with three NEG-based probes on all tasks in MCEval8K. For other LLMs, to reduce computational cost, we only probe one dataset per genre (NER, Agnews, PAWS, CSQA, HaluEval, and mLAMA) with $2^{10}$ training examples, focusing on majority-vote probes.
@@ 5.5 Result and Analysis
We compare our skill probes with three baselines: Rand (random guessing, LM-Prob (choosing the token with highest model probability), and Act (activation-based probes from Wang et al. (2022)).
Can NEG capture language skills?
Figure 6 shows Llama2-7B’s accuracy on MCEval8K tasks in zero- and few-shot settings. The results show that LM-Prob consistently outperforms random guessing, indicating that Llama2-7B can follow instructions to activate the relevant language skills. However, even the simple majority-vote probes can surpass the LM-Prob, suggesting that NEGs capture meaningful information about language skills in individual neurons. Furthermore, Tree-Probe significantly outperforms majority-vote probes by integrating NEGs from multiple neurons. This highlights that the interplay among neurons is crucial for accurately representing language skills. See Tables 8 and 9 in Appendix § G.1 for details on all tasks. Experiments on other LLMs further validate our findings as reported in Table 3.
Which encodes skills better: activation or gradient?
Table 3 demonstrates that NEG-based probes have a different focus from the activation-based approach. Specifically, the gradient-based probes perform better on knowledge-intensive language tasks, such as factual knowledge and commonsense reasoning (mLAMA and CSQA). One possible explanation is that simple linguistic knowledge becomes saturated during pretraining, while complex knowledge remains undertrained, making it more effectively captured by NEG-based probes.
Furthermore, NEG-based probes outperform activation-based methods in efficiency by assigning distinct values to different target tokens per neuron, enhancing representational capacity. In contrast, activation-based approaches rely only on forward pass signals, limiting neurons to binary distinctions and making them less effective for multi-class tasks Wang et al. (2022).
How do NEGs make predictions?
We analyze NEGs associated with each option token to explore why simple majority-vote classifiers can achieve high accuracy. Using PAWS (binary classification) as an example, we acquire NEGs for target tokens (“yes/no”) across 6K prompts. We find 97.21% of neurons show opposite signs for “yes/no” tokens, with a near-perfect negative correlation ($-0.9996$), indicating strong polarity across options. We further compare the NEGs given zero-shot and few-shot prompts. We find that NEG magnitudes of few-shot prompts over 22 tasks are 5.36 times larger than zero-shot, suggesting that adding demonstrations can effectively activate skill neurons.
6 Properties of Skill Neurons
To deepen our understanding of how language skills are represented within neurons, we analyze skill neurons’ behavioral and structural properties through a series of focused studies.
Neuron sizes Tasks
$2^{0}\sim 2^{3}$ Toxic, LTI, M-POS, FEVER, TempLAMA
$2^{9}\sim 2^{13}$ Agnews, MyriadLAMA, CSQA, mLAMA
Table 4: The optimal neuron sizes for Mag-Probe.
Figure 7: Accuracies with varying training sizes.
6.1 Representation & Acquisition Efficiency
Representational efficiency:
By tuning neuron size on the validation set, we find that skill-neuron probes achieve high accuracy with only a few neurons. The optimal neuron sizes for all tasks with Magn-Probe are reported in Table 4. Most tasks can achieve optimal accuracy within 256 neurons, demonstrating the efficiency of NEG in representing language skills. Notably, factuality tasks, such as MyriadLAMA, CSQA, and mLAMA, engage more neurons, suggesting that handling facts requires more diverse neurons, reflecting the complexity of factual understanding tasks.
Acquisition efficiency:
Figure 7 reports the accuracy of skill-neuron probes with different training examples. While adding training examples can consistently increase the probes’ accuracy, the earnings slow down after 128, indicating the efficiency of acquiring skill neurons with limited data.
6.2 How Robust Are Skill Neurons to Contextual Variations?
We investigate how skill neurons change when we provide different contexts, including instructions, demonstrations, and options for the same task. Given context $X$, we first acquire the skill neurons $\mathcal{N}_{s}^{X}$ and the accuracy $\text{ACC}_{\mathcal{N}_{s}^{X}}^{X}$. Then, we use the classifier built with $\mathcal{N}_{s}^{X}$ to evaluate the task by context $Y$ as $\text{ACC}_{\mathcal{N}_{s}^{X}}^{Y}$. We denote the robustness of $\mathcal{N}_{s}^{X}$ on context $Y$ as $\frac{\max(\text{ACC}_{\mathcal{N}_{s}^{X}}^{Y}-\alpha,0)}{\max(\text{ACC}_{\mathcal{N}_{s}^{Y}}^{Y}-\alpha,0)}$, where $\alpha$ is the accuracy by Rand.
Using PAWS as an example, we create 12 distinct contexts by varying the instructions, the selection of demonstrations, and the output token styles. By measuring robustness across combinations, we find that it remains very high (near 1) for varying instructions and demonstrations but drops significantly when target tokens change. The results show that skill neurons remain highly robust to changes in instructions and demonstrations but lose robustness when output tokens change. See § J for details of experimental settings and results.
6.3 Are Neurons Substitutable in Representing Language Skills?
We investigate whether skill neurons uniquely represent skills or can be substituted by different neurons. For investigation, we build Magn-Probes using different neuron sets. Specifically, we select 64 consecutive neurons from the ranked list, ordered by their importance as skill indicators (§ 5.3). [^12] We use 64-neuron units that maintain high accuracy across tasks (§ A). With 352,256 neurons in Llama2-7B’s FF layers, this produces 5,504 unique neuron sets.
Figure 8 depicts the accuracies across six tasks with different neuron sets, showing that representational accuracy declines only gradually with less important neurons. Even using the least important neurons still outperforms Rand. This suggests that numerous neurons can act as skill indicators and skill neurons are broadly distributed and substitutable (see § H.2 for full results).
Figure 8: Accuracies of Magn-Probe with different neuron sets, plotting the mean accuracy within each window, along with the accuracy ranges, as the envelope. Neuron sets are selected from all neurons in Llama2-7B in groups of 32, ranked by importance as skill indicators.
Figure 9: Accuracies of Tree-Probes with varying depths and trees. X-axis: logarithm of trees’ number; Y-axis: logarithm of tree depths; Z-axis: Accuracy.
6.4 Do Skill Neurons Depend on Each Other?
The majority-vote probes assume neuron independence, while the Tree-Probe model their interdependencies via hierarchical classifiers. Figure 6 shows that Tree-Probles better represents language skills by accounting for neuron interdependency. To see how important interdependency is in encoding language skills, we train Tree-Probes with varying #n_trees and #n_layers.131313#n_trees and #n_layers are set to $2^{N}$ and $2^{M}$, where $0\leq N\leq 10$, $1\leq M\leq 11$, and $N+M<12$.
Figure 9 reports the resulting accuracies of Tree-Probe on PAWS, CSQA, and HaluEval. Their different shapes indicate the interdependency style between neurons for diverse skills differs. Some tasks (PAWS) prefer deep layers, while some (CSQA) prefer more trees, and some (HaluEval) require a balance. See § H.3 for details.
7 Related Work
Neuron-level knowledge attribution methods. Previous work links neurons to knowledge Oba et al. (2021) by measuring their impact on model predictions. Some use causal interventions Geva et al. (2021); Meng et al. (2022); Chen et al. (2023); Wang et al. (2024), while others rely on heavy tensor calculations Geva et al. (2022); Lee et al. (2024); Yu and Ananiadou (2024). These methods are computationally costly, limiting their use for large-scale probing across diverse LLM prompts. They also focus on relative neuron rankings rather than the precise, quantitative neuron-output relationships, reducing their applicability in use cases like knowledge editing Zhang et al. (2024) and bias mitigation Gallegos et al. (2024). Gradient-based approaches (Lundstrom et al., 2022; Dai et al., 2022) similarly face high computational costs.
Skill neuron probing. Neurons in feed-forward layers can encode specific language skills, enabling the tasks to be solved using only their activations; such neurons are called skill neurons Wang et al. (2022); Song et al. (2024). Prior work showed that neurons represent semantic skills like sentiment classification Wang et al. (2022); Song et al. (2024) and complex skills such as style transfer Lai et al. (2024) and translation Tan et al. (2024). However, these studies focus on activations as knowledge indicators, overlooking neuron gradients’ potential to represent language skills, which limits neuron-level model adjustments.